OpenStudy (anonymous):

prove (cos4x+cos3x)/(sin4x-sin3x) = cotx/2 Can anyone help me explain the answer please

OpenStudy (anonymous):

cos4x + cos3x = 2cos(7x/2)cos(x/2) sin4x - sin3x = 2cos(7x/2)sin(x/2) Their division is given. On dividing cos(7x/2) Cancels out. We are left with. cos(x/2) / sin(x/2) = cot(x/2)

OpenStudy (anonymous):

@siddhantsharan you typed fast!!!

OpenStudy (anonymous):

HAha. Thanks. :D

OpenStudy (anonymous):

how did you get 2cos(7x/2)cos(x/2)

OpenStudy (anonymous):

someone help me pleaseeee

OpenStudy (anonymous):

are the 4 and 3 meant to be exponents?

OpenStudy (anonymous):

nooo

OpenStudy (anonymous):

Then it's not an identity, you can't prove it.

OpenStudy (anonymous):

but i got the first answer

OpenStudy (anonymous):

I just wolfram alphad it, it's not an identity.

OpenStudy (anonymous):

how do you do it then?

OpenStudy (dumbcow):

http://www.mathwords.com/s/sum_to_product_identities.htm the 1st post has it correct, here is how they came up with it

OpenStudy (anonymous):

you're saying wolfram alpha is wrongggggggggg?

OpenStudy (dumbcow):

i am, haha post the link

OpenStudy (anonymous):

OMG thank you! i dont have that pellet on my identity sheet

OpenStudy (dumbcow):

welcome, yeah i never learned these identities very well either